# These $\equalcontThese authors contributed equally to this work.$

### v2

@@ 5.4.1 Comparison Sampling Strategies
Tables 6, 7 present data to compare metrics and computational resource usage for training with $\mathrm{CCE}^{-}$ using global uniform and popularity-based sampling strategies. In terms of $\mathrm{NDSG@10}$ metric, global uniform sampling consistently outperforms the popularity-based approach across all considered datasets. However, popularity-based sampling shows significant gains in diversity-related metrics for the Gowalla and Beauty datasets. Compared to popularity-based sampling, using global uniform sampling leads to a reduction in $\mathrm{Coverage@10}$ by 185.8% and 63.9 %, and a reduction in $\mathrm{Surprisal@10}$ by 40% and 19.8%. These results align with observations that popularity-weighted sampling enhances exposure to long-tail items, improving recommendation diversity and novelty [35].
@@
Another approach to negative sampling is the Scalable Cross-Entropy (SCE) method  [21]. This method approximates the CE loss by employing a sampling procedure that computes logits only for elements with the largest absolute values, rather than for all items in the catalog. As a result, memory usage for the logit tensor is significantly reduced, and the loss computation is accelerated.
@@
Table 8: Results for the SCE loss with the SASRec model
@@
(bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
@@
Table 9: Results for the comparison of SCE loss with $\mathrm{CCE}^{-}$ for the SASRec model. $\Delta$: Relative difference compared to SCE and $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
Table 10: Results for the comparison of SCE loss with CCE for the SASRec model. $\Delta$: Relative difference compared to SCE and $\mathrm{CCE}^{-}$, green denotes CCE superiority.
@@
5.5 Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE
As outlined in the original paper [15], the SASRec model can be trained using the Binary Cross-Entropy (BCE) loss. To evaluate the effectiveness of our proposed methods, we conducted training SASRec with the BCE loss.
@@ 5.5 Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE
Table 11: Results for the BCE loss with the SASRec model.
@@
(bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
@@
Table 12: Results for the comparison of BCE loss with $\mathrm{CCE}^{-}$ for the SASRec model. $\Delta$: Relative difference compared to BCE and $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
Table 13: Results for the comparison of BCE loss with CCE for the SASRec model. $\Delta$: Relative difference compared to BCE and CCE, green denotes CCE superiority.
@@
5.6 Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
In addition to the SASRec architecture, we study the efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec  [29].
@@ 5.6 Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
Table 14: Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using identical training hyperparameters. Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(bs/sl/ns) NDCG Mem, Time, NDCG Mem, Time, % % %
@@
Table 15: Comparison Coverage@10 and Surprisal@10 for CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using identical hyperparameters. Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(bs/sl/ns) Coverage Surprisal Coverage Surprisal % %
@@
Table 16: Performance of Bert4Rec with CE loss.
@@
(bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
@@
Table 17: Performance of Bert4Rec with $\mathrm{CE}^{-}$ loss.
@@
(bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s

### v3

@@ 5.4.1. Comparison Sampling Strategies
Tables 6, 7 present data to compare metrics and computational resource usage for training with $\mathrm{CCE}^{-}$ using global uniform and popularity-based sampling strategies. In terms of $\mathrm{NDSG@10}$ metric, global uniform sampling consistently outperforms the popularity-based approach across all considered datasets. However, popularity-based sampling shows significant gains in diversity-related metrics for the Gowalla and Beauty datasets. Compared to popularity-based sampling, using global uniform sampling leads to a reduction in $\mathrm{Coverage@10}$ by 185.8% and 63.9 %, and a reduction in $\mathrm{Surprisal@10}$ by 40% and 19.8%. These results align with observations that popularity-weighted sampling enhances exposure to long-tail items, improving recommendation diversity and novelty (wang2024popularitylearning).
@@
Another approach to negative sampling is the Scalable Cross-Entropy (SCE) method  (mezentsev2024scalable). This method approximates the CE loss by employing a sampling procedure that computes logits only for elements with the largest absolute values, rather than for all items in the catalog. As a result, memory usage for the logit tensor is significantly reduced, and the loss computation is accelerated.
@@
Table 8. Results for the SCE loss with the SASRec model
@@
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
@@
Table 9. Results for the comparison of SCE loss with $\mathrm{CCE}^{-}$ for the SASRec model. $\Delta$: Relative difference compared to SCE and $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
Table 10. Results for the comparison of SCE loss with CCE for the SASRec model. $\Delta$: Relative difference compared to SCE and $\mathrm{CCE}^{-}$, green denotes CCE superiority.
@@
$5.$ Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE
As outlined in the original paper (kang2018self), the SASRec model can be trained using the Binary Cross-Entropy (BCE) loss. To evaluate the effectiveness of our proposed methods, we conducted training SASRec with the BCE loss.
@@
Table 11. Results for the BCE loss with the SASRec model.
@@
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
@@
Table 12. Results for the comparison of BCE loss with $\mathrm{CCE}^{-}$ for the SASRec model. $\Delta$: Relative difference compared to BCE and $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
Table 13. Results for the comparison of BCE loss with CCE for the SASRec model. $\Delta$: Relative difference compared to BCE and CCE, green denotes CCE superiority.
@@
5.6. Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
In addition to the SASRec architecture, we study the efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec  (sun2019bert4rec).
@@ 5.6. Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
Table 14. Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using identical training hyperparameters. Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(lr)3-5 (lr)6-8 (bs/sl/ns) NDCG Mem, Time, NDCG Mem, Time, % % %
@@
Table 15. Comparison Coverage@10 and Surprisal@10 for CCE and $\mathrm{CCE}^{-}$ on the Bert4Rec model using identical hyperparameters. Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(lr)3-4 (lr)5-6 (bs/sl/ns) Coverage Surprisal Coverage Surprisal % %
@@
Table 16. Performance of Bert4Rec with CE loss.
@@
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s
@@
Table 17. Performance of Bert4Rec with $\mathrm{CE}^{-}$ loss.
@@
(lr)3-7 (bs/sl) NDCG@10 Coverage@10 Surprisal@10 Mem, GB Time, s